(a-5)^2+(b-c)^2+(c-d)^2+(b+c+d-9)^2=0 then (a+b+c)(b+c+d) ?

  1. (a-5)^2+(b-c)^2+(c-d)^2+(b+c+d-9)^2=0 then (a+b+c)(b+c+d) ?

Grade:12

1 Answers

Subhiksha
8 years ago
If x^2 + y^2 + z^2 =0, then x=y=z=0.
 
Here,
a=5
b=c=d
3b=9 or b=3=c=d
 
so, substituting these values
(a+b+c)(b+c+d)= 11*9=99